# Calogero–Sutherland hyperbolic system and Heckman–Opdam $\mathfrak{gl}_{n}$ hypergeometric function

### v1

@@ 5.1 Asymptotics and bounds
For $g>1$, ${\bm{\lambda}}_{n}\in\mathbb{R}^{n}$ and arbitrary small $\varepsilon>0$ there exist constants $N(g),C(g,\varepsilon)>0$ such that in the region (5.5)
@@ 5.2 Duality
$\begin{split}\mu(\bm{y}_{n-1})=w(\bm{y}_{n-1})w(-\bm{y}_{n-1}),\qquad\hat{\mu}(\bm{y}_{n-1})=\hat{w}(\bm{y}_{n-1})\hat{w}(-\bm{y}_{n-1}),\end{split}$ (5.16)
@@ 6 Heckman–Opdam hypergeometric function
In the previous sections we considered wave functions of Calogero–Sutherland model constructed using integral operators. Alternatively, Heckman and Opdam studied wave functions expressed in terms of so-called Harish-Chandra series [HO]. In particular, they proved that Calogero–Sutherland Hamiltonians admit unique eigenfunctions $F_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$ with normalization
@@
${\Psi}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})={\hat{\Psi}}_{{\bm{\lambda}}_{n}}(\bm{x}_{n}).$
@@
${\Psi}^{as}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})=e^{-2\pi g(\bm{x}_{n},\bm{\rho}_{n})}\sum_{\sigma\in S_{n}}e^{2\pi\imath\sum_{i=1}^{n}x_{i}\lambda_{\sigma(i)}}\prod_{{i<j}}\hat{w}^{-1}(\lambda_{\sigma(i)}-\lambda_{\sigma(j)}).$ (6.26)
Theorem 5.1 implies the following asymptotical behaviour of the function ${\Psi}^{as}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$ for real ${\bm{\lambda}}_{n}$.
@@
Since $\Psi_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$ is symmetric in $\bm{x}_{n}$, in the region
@@
On the other hand, the asymptotics $F^{as}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$ of the Heckman–Opdam $A_{n-1}$ hypergeometric function $F_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$ (extended to $\mathfrak{gl}_{n}$) in the region (6.28) is given by [HO, O]
@@
In Appendix C we also verify by direct calculation that the normalization condition (6.25) is satisfied.
@@
$\hat{M}_{r}\,\prod_{i\not=j}\Gamma(\imath(\lambda_{i}-\lambda_{j})+g)=\prod_{i\not=j}\Gamma(\imath(\lambda_{i}-\lambda_{j})+g)\,\hat{H}_{r}.$
@@
${F}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$ $=\alpha_{n}\prod_{i\not=j}\Gamma^{-1}(\imath(\lambda_{i}-\lambda_{j})+g)\ \Lambda_{n}(\lambda_{n})\,\Lambda_{n-1}(\lambda_{n-1})\cdots\Lambda_{2}(\lambda_{2})\,e^{{2\pi\imath}\lambda_{1}x_{1}}$ (6.38)
$=\alpha_{n}\prod_{i\not=j}\Gamma^{-1}(\imath(\lambda_{i}-\lambda_{j})+g)\ \hat{\Lambda}_{n}(x_{n})\,\hat{\Lambda}_{n-1}(x_{n-1})\cdots\,{\hat{\Lambda}}(x_{2})\,e^{2\pi i\lambda_{1}x_{1}}.$ (6.39)
Standard bounds on gamma function $\Gamma(\lambda)$ (Stirling formula in the cone $|\arg\lambda|<\pi$ together with reflection formula between the poles $\lambda\in\mathbb{Z}_{\leq 0}$) show that Mellin–Barnes integral (4.90) can be calculated by residue technique in the region $x_{1}>x_{2}>\ldots>x_{n}$ with $m_{n}(x)>\delta$ for some fixed $\delta>0$. The result is the convergent Harish-Chandra series, which we desribe now.
@@
The residue calculation shows that the function $\Psi_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$ is the symmetrization over spectral variables of the function $\psi_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$.
@@
This formula can be regarded as a rational counterpart of the Noumi–Shiraishi series [NS].
@@ Acknowledgments
We are immensely grateful to S. Kharchev for collaboration on the previous stage of the present work and M. Hallnäs for useful discussions. Sections 3 and 4 of this work were done within a research project implemented as part of the Basic Research Program at the HSE University. Section 2 and Appendix B were written with the support of the Grant from the President of the Russian Federation for students of Master’s degree programs.
@@ A.1 Double sine function
$\hat{w}^{R}(\lambda)=S_{2}(\imath\lambda|\hat{\bm{\omega}})\,S_{2}\Big{(}-\imath\lambda+\frac{\mathrm{g}}{\omega_{1}\omega_{2}}\,\Big{|}\hat{\bm{\omega}}\Big{)}$, (A.57)
$\hat{K}^{R}(\lambda)=S_{2}^{-1}\Bigl{(}\imath\lambda+\frac{\mathrm{g}}{2\omega_{1}\omega_{2}}\Big{|}\hat{\bm{\omega}}\Bigr{)}\,S_{2}^{-1}\Bigl{(}-\imath\lambda+\frac{\mathrm{g}}{2\omega_{1}\omega_{2}}\Big{|}\hat{\bm{\omega}}\Bigr{)}$. (A.58)
@@ A.2 Uniform limits of kernel and measure functions
$\begin{split}\Pi_{1}(\varepsilon)=\{z\colon\leavevmode\nobreak\ \varepsilon&\leq\mathrm{Re}\,(z)\leq 1-\varepsilon\},\\[6.0pt] \Pi_{2}(\varepsilon)=\{z\in\imath\mathbb{R}\colon\leavevmode\nobreak\ |\operatorname{Im}(z)|\geq\varepsilon\},&\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \Pi_{3}(\varepsilon)=\{z\in\imath\mathbb{R}\colon\leavevmode\nobreak\ |\operatorname{Im}(z)|<\varepsilon\}.\end{split}$ (A.61)
@@
$\left|\frac{S_{2}(z)}{S_{2}(z+g\omega_{2})}\right|=\big{|}\big{(}2\sin(\pi z)\big{)}^{g}\big{|}(1+O(\omega_{2})),\leavevmode\nobreak\ \leavevmode\nobreak\ z\in\Pi_{1}(\varepsilon),\leavevmode\nobreak\ \leavevmode\nobreak\ 0\leq\omega_{2}\leq a(\varepsilon).$ (A.62)
@@
$\left|\frac{S_{2}(z)}{S_{2}(z+g\omega_{2})}\right|=\big{|}\big{(}2\sin(\pi z)\big{)}^{g}\big{|}(1+O(\omega_{2})),\leavevmode\nobreak\ \leavevmode\nobreak\ z\in\Pi_{2}(\varepsilon),\leavevmode\nobreak\ \leavevmode\nobreak\ 0\leq\omega_{2}\leq a(\varepsilon).$ (A.63)
@@
$\left|\frac{S_{2}(z)}{S_{2}(z+g\omega_{2})}\right|\leq C(\varepsilon),\leavevmode\nobreak\ \leavevmode\nobreak\ z\in\Pi_{3}(\varepsilon),\leavevmode\nobreak\ \leavevmode\nobreak\ 0\leq\omega_{2}\leq a(\varepsilon).$ (A.64)
@@
$\left|K^{R}(x|1,\omega_{2},g\omega_{2})\right|=\left|\frac{S_{2}(\imath x+\frac{1+\omega_{2}}{2}+\frac{g\omega_{2}}{2})}{S_{2}(\imath x+\frac{1+\omega_{2}}{2}-\frac{g\omega_{2}}{2})}\right|=\left|(2\operatorname{ch}(\pi x))^{-g}\right|\left(1+O(\omega_{2})\right).$ (A.65)
@@
$\left|\mu^{R}(x|1,\omega_{2},g\omega_{2})\right|=\left|\frac{S_{2}(\imath x)}{S_{2}(\imath x+g\omega_{2})}\frac{S_{2}(-\imath x)}{S_{2}(-\imath x+g\omega_{2})}\right|=\left|\left(4\operatorname{sh}^{2}(\pi x)\right)^{g}\right|\left(1+O(\omega_{2})\right),$ (A.66)
@@
$\log\frac{S_{2}(z+g\omega_{2})}{S_{2}(z)}=\frac{\pi\imath g}{2}\left({2z}+{(g-1)\omega_{2}}-1\right)+\int_{\mathbb{R}+\imath 0}\frac{e^{zt}(e^{g\omega_{2}t}-1)}{(e^{t}-1)(e^{\omega_{2}t}-1)}\frac{dt}{t}.$ (A.69)
@@
$\begin{split}&\log\frac{S_{2}(z)}{S_{2}(z+g\omega_{2})}-g\log\big{(}2\sin(\pi z)\big{)}=-\frac{\pi\imath\omega_{2}}{2}g(g-1)\\[6.0pt] &-\int_{\mathbb{R}+\imath 0}\frac{e^{zt}}{e^{t}-1}\left(\frac{e^{g\omega_{2}t}-1}{e^{\omega_{2}t}-1}-g\right)\frac{dt}{t}=-\,v.p.\int_{\mathbb{R}}\frac{e^{zt}}{e^{t}-1}\left(\frac{e^{g\omega_{2}t}-1}{e^{\omega_{2}t}-1}-g\right)\frac{dt}{t}\end{split}$ (A.70)
@@
$\log\big{|}2\sin(\pi z)\big{|}=\pi\operatorname{Im}z-\mathrm{Re}\,\int_{\mathbb{R}+\imath 0}\frac{e^{zt}}{e^{t}-1}\frac{dt}{t},\qquad 0\leq\mathrm{Re}\,(z)<1,\qquad z\not=0,$ (A.71)
@@
$\hskip 199.16928pt0\leq\mathrm{Re}\,(z)<1+(1-g)\omega_{2},\qquad z\not=0.$
@@
Proofs of Propositions A.1A.3. Using Lemmas A.1A.2 and A.3 we obtain
$\begin{split}\left|\mathrm{Re}\,\left(\log\frac{S_{2}(z)}{S_{2}(z+g\omega_{2})}-g\log\big{(}2\sin(\pi z)\big{)}\right)\right|\leq\omega_{2}(C_{1}+C_{2}+C_{3}),\\ 0\leq\omega_{2}\leq a(\varepsilon),\leavevmode\nobreak\ \leavevmode\nobreak\ z\in\Pi_{1}(\varepsilon).\end{split}$ (A.79)
@@
$\begin{split}e^{-\omega_{2}(C_{1}+C_{2}+C_{3})}\leq\left|\frac{S_{2}(z)}{S_{2}(z+g\omega_{2})}\cdot\big{(}2\sin(\pi z)\big{)}^{-g}\right|\leq e^{\omega_{2}(C_{1}+C_{2}+C_{3})},\\ 0\leq\omega_{2}\leq a(\varepsilon),\leavevmode\nobreak\ \leavevmode\nobreak\ z\in\Pi_{1}(\varepsilon),\end{split}$ (A.80)
@@
To prove Proposition A.3 apply Lemmas A.1A.2 and A.5 to (LABEL:closed)

### v3

@@ 5.1 Asymptotics and bounds
For $g>1$, ${\bm{\lambda}}_{n}\in\mathbb{R}^{n}$ and $\varepsilon\in(0,2\pi)$ there exist constants $N(g),C(g,\varepsilon)>0$ such that in the region (5.5)
@@ 5.2 Duality
$\begin{split}\mu(\bm{y}_{n-1})=w(\bm{y}_{n-1})w(-\bm{y}_{n-1}),\qquad\hat{\mu}(\bm{y}_{n-1})=\hat{w}(\bm{y}_{n-1})\hat{w}(-\bm{y}_{n-1}),\end{split}$
@@ 6 Heckman–Opdam hypergeometric function
In the previous sections we considered wave functions of Calogero–Sutherland model constructed using integral operators. Alternatively, Heckman and Opdam studied wave functions expressed in terms of so-called Harish-Chandra series [HO]. In particular, they proved that Calogero–Sutherland Hamiltonians (3.3) admit unique eigenfunctions $F_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$ with normalization
@@
${\Psi}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})={\hat{\Psi}}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})=\frac{1}{\hat{w}({\bm{\lambda}}_{n})}\,\Phi_{{\bm{\lambda}}_{n}}(\bm{x}_{n}).$
@@
${\Psi}^{as}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})=\frac{1}{\hat{w}({\bm{\lambda}}_{n})}\,\Phi^{as}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})=e^{-2\pi g(\bm{x}_{n},\bm{\rho}_{n})}\sum_{\sigma\in S_{n}}e^{2\pi\imath\sum_{i=1}^{n}x_{i}\lambda_{\sigma(i)}}\prod_{{i<j}}\hat{w}^{-1}(\lambda_{\sigma(i)}-\lambda_{\sigma(j)}).$ (6.2)
Theorem 5.1 implies the following asymptotical behaviour of the function ${\Psi}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$ for real ${\bm{\lambda}}_{n}$.
@@
Here the factor $\hat{\Delta}({\bm{\lambda}}_{n})^{N(g)}$ from Theorem 5.1 is hidden in the constant $C(g,{\bm{\lambda}}_{n})$. Since $\Psi_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$ is symmetric in $\bm{x}_{n}$, in the region
@@
On the other hand, the asymptotics $F^{as}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$ of the Heckman–Opdam $A_{n-1}$ hypergeometric function $F_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$ (extended to $\mathfrak{gl}_{n}$) in the region (6.4) is given by [HO, O1]
@@
In Appendix C we also prove the following proposition by calculating the Mellin–Barnes integral (4.14) at the zero point, which verifies the normalization condition (6.1).
Proposition 6.2.
For ${\bm{\lambda}}_{n}\in\mathbb{R}^{n}$ we have
$\Psi_{\bm{\lambda}_{n}}(0,\dots,0)=\prod_{k=1}^{n}\frac{\Gamma(g)}{\Gamma(kg)}\,\prod_{1\leq i\not=j\leq n}\,\Gamma(\imath(\lambda_{i}-\lambda_{j})+g).$ (6.11)
@@
$\hat{M}_{r}\,\prod_{i\not=j}\Gamma(\imath(\lambda_{i}-\lambda_{j})+g)=\prod_{i\not=j}\Gamma(\imath(\lambda_{i}-\lambda_{j})+g)\,\hat{H}_{r}.$ (6.14)
@@
${F}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$ $=\alpha_{n}\prod_{i\not=j}\Gamma^{-1}(\imath(\lambda_{i}-\lambda_{j})+g)\ \Lambda_{n}(\lambda_{n})\,\cdots\Lambda_{2}(\lambda_{2})\,e^{{2\pi\imath}\lambda_{1}x_{1}}$
$=\alpha_{n}\prod_{i\not=j}\Gamma^{-1}(\imath(\lambda_{i}-\lambda_{j})+g)\ \hat{\Lambda}_{n}(x_{n})\,\cdots\,{\hat{\Lambda}}(x_{2})\,e^{2\pi i\lambda_{1}x_{1}}.$
Conversely, the Mellin–Barnes integrals are standardly converted into series by residue technique. For example, in the simplest case $n=2$ we have
$\Psi_{\lambda_{1},\lambda_{2}}(x_{1},x_{2})=\frac{1}{2\pi\Gamma(g)}\int_{\mathbb{R}}d\gamma\;e^{2\pi\imath x_{2}(\lambda_{1}+\lambda_{2})+2\pi\imath(x_{1}-x_{2})\gamma}\\[6.0pt] \times\prod_{j=1,2}\Gamma\biggl(\imath(\lambda_{j}-\gamma)+\frac{g}{2}\biggr)\,\Gamma\biggl(\imath(\gamma-\lambda_{j})+\frac{g}{2}\biggr),$ (6.16)
which for $x_{1}>x_{2}$ can be evaluated by encircling the contour in the upper half-plane. Two series of the integrand poles
$\gamma=\lambda_{j}+\imath\frac{g}{2}+\imath m,\qquad m\in\mathbb{Z}_{\geq 0},\qquad j=1,2$ (6.17)
lead to the splitting of the wave function
$\Psi_{\lambda_{1},\lambda_{2}}(x_{1},x_{2})=\psi_{\lambda_{1},\lambda_{2}}(x_{1},x_{2})+\psi_{\lambda_{2},\lambda_{1}}(x_{1},x_{2})$ (6.18)
into two series of hypergeometric type
$\psi_{\lambda_{1},\lambda_{2}}(x_{1},x_{2})=e^{2\pi\bigl(\imath x_{1}\lambda_{1}+\imath x_{2}\lambda_{2}-\frac{g}{2}(x_{1}-x_{2})\bigr)}\sum_{m=0}^{\infty}\frac{(-1)^{m}\,\Gamma(g+m)}{m!\,\Gamma(g)}\\[6.0pt] \times\Gamma(\imath(\lambda_{1}-\lambda_{2})-m)\,\Gamma(\imath(\lambda_{2}-\lambda_{1})+m+g)\,e^{2\pi m(x_{2}-x_{1})}.$ (6.19)
As explained in [KK1, Section 4], the formula (6.18) is equivalent to the relation between associated Legendre functions of the first and the second kinds [DLMF, (14.9.12)]
$\frac{e^{-\pi\imath\mu}}{\cos(\pi\nu)}\bigl(Q_{-\nu-1}^{\mu}(z)-Q_{\nu}^{\mu}(z)\bigr)=\Gamma(\mu+\nu+1)\,\Gamma(\mu-\nu)\,P_{\nu}^{-\mu}(z).$ (6.20)
Below we describe its generalization to the case $n>2$.
@@
This is the Harish-Chandra type series generalizing the two-variable functions (6.19). The residue calculation shows that the wave function $\Psi_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$ is the symmetrization over spectral variables of the function $\psi_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$.
@@
Proof.
The proof is by induction. The base case $n=1$ is trivial. For $n>1$ we evaluate the Mellin–Barnes representation (4.14)
$\Psi_{{\bm{\lambda}}_{n}}(\bm{x}_{n})=\hat{d}_{n-1}(g)\int_{\mathbb{R}^{n-1}}d\bm{\gamma}_{n-1}\;\hat{\Lambda}_{n-1}({\bm{\lambda}}_{n},\bm{\gamma}_{n-1};x_{n})\,\Psi_{\bm{\gamma}_{n-1}}(\bm{x}_{n-1})$ (6.26)
by residue technique. The integrand has poles in the upper half-plane at the points
$\gamma_{j}=\lambda_{\sigma(j)}+\imath\frac{g}{2}+\imath m_{j},\qquad m_{j}\in\mathbb{Z}_{\geq 0},\qquad\sigma\in S_{n}.$ (6.27)
Notice that there are no poles with one $\lambda_{i}$ for two integration variables $\gamma_{j}$ and $\gamma_{k}$, since the measure function
$\hat{\mu}(\bm{\gamma}_{n-1})=\prod_{j\neq k}\Gamma^{-1}(\imath(\gamma_{j}-\gamma_{k}))\,\Gamma^{-1}(\imath(\gamma_{j}-\gamma_{k})+g)$ (6.28)
has zeros at the points $\gamma_{j}-\gamma_{k}\in\imath\mathbb{Z}$. Besides, by Theorem 6.1 the function
$\prod_{j\neq k}\Gamma^{-1}(\imath(\gamma_{j}-\gamma_{k})+g)\,\Psi_{\bm{\gamma}_{n-1}}(\bm{x}_{n-1})$
modulo constant coincides with the Heckman–Opdam function $F_{\bm{\gamma}_{n-1}}(\bm{x}_{n-1})$, which is known to be entire in spectral parameters [O2, Theorem 2.8].
Summing up residues at the poles (6.27), using induction assumption and symmetry of the kernel $\hat{\Lambda}_{n-1}({\bm{\lambda}}_{n},\bm{\gamma}_{n-1};x_{n})$ with respect to $\gamma_{j}$ we obtain the expansion
$\Psi_{{\bm{\lambda}}_{n}}(\bm{x}_{n})=\sum_{\sigma\in S_{n}}\tilde{\psi}_{\lambda_{\sigma(1)},\ldots,\lambda_{\sigma(n)}}(\bm{x}_{n})$ (6.29)
in terms of the series
$\tilde{\psi}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})=\sum_{m_{1},\dots,m_{n-1}=0}^{\infty}A_{\bm{m}_{n-1}}({\bm{\lambda}}_{n};x_{n})\,\psi_{{\bm{\lambda}}_{n-1}+\imath\frac{g}{2}\bm{e}_{n-1}+\imath\bm{m}_{n-1}}(\bm{x}_{n-1}),$ (6.30)
where $\bm{e}_{n-1}=(1,\dots,1)$ and the coefficients
$A_{\bm{m}_{n-1}}({\bm{\lambda}}_{n};x_{n})=(n-1)!\,\hat{d}_{n-1}(g)\Biggl[\prod_{j=1}^{n-1}2\pi\imath\operatorname{Res}_{\gamma_{j}=\lambda_{j}+\imath\frac{g}{2}+\imath m_{j}}\Biggr]\,\hat{\Lambda}({\bm{\lambda}}_{n},\bm{\gamma}_{n-1};x_{n}).$ (6.31)
Explicitly, they are given by
$A_{\bm{m}_{n-1}}({\bm{\lambda}}_{n})=e^{2\pi\imath x_{n}\bigl(\lambda_{n}-\imath g\frac{n-1}{2}-\imath\underline{\bm{m}}_{n-1}\bigr)}\,\prod_{j=1}^{n-1}\frac{(-1)^{m_{j}}\,\Gamma(g+m_{j})}{m_{j}!\,\Gamma(g)}\\ \times\frac{\prod\limits_{s=1}^{n-1}\prod\limits_{\begin{subarray}{c}t=1\\ t\neq s\end{subarray}}^{n}\Gamma(\imath(\lambda_{s}-\lambda_{t})-m_{s})\,\Gamma(\imath(\lambda_{t}-\lambda_{s})+m_{s}+g)}{\prod\limits_{1\leq s\neq t\leq n-1}\Gamma(\imath(\lambda_{s}-\lambda_{t})-m_{s}+m_{t})\,\Gamma(\imath(\lambda_{t}-\lambda_{s})+m_{s}-m_{t}+g)}.$ (6.32)
It is straightforward to check that $\tilde{\psi}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})=\psi_{{\bm{\lambda}}_{n}}(\bm{x}_{n})$.
Standard bounds on gamma function $\Gamma(\lambda)$ (Stirling formula in the cone $|\arg\lambda|<\pi$ together with reflection formula between the poles $\lambda\in\mathbb{Z}_{\leq 0}$) show that the above calculation is justified in the region $x_{1}>x_{2}>\ldots>x_{n}$ with $m_{n}(x)>\delta$ for some fixed $\delta>0$. ∎
@@
The formula (6.24) can be regarded as a rational counterpart of the Noumi–Shiraishi series [NS].
@@ Acknowledgments
We are immensely grateful to S. Kharchev for collaboration on the previous stage of the present work and M. Hallnäs for useful discussions. S. D. and S. Kh. thank BIMSA for hospitality. A big part of the work was done during their visit to BIMSA. Section 3 of this work was done within a research project implemented as part of the Basic Research Program at the HSE University. Section 4 was written under support by Russian Science Foundation, project No. 23-11-00150. Section 2 and Appendix B were written with the support of the Grant from the President of the Russian Federation for students of Master’s degree programs.
@@ A.1 Double sine function
[…]
@@ A.2 Uniform limits of kernel and measure functions
$\begin{split}\Pi_{1}(\varepsilon)=&\{z\colon\penalty 10000\ \varepsilon\leq\mathrm{Re}\,(z)\leq 1-\varepsilon\},\\[6.0pt] \Pi_{2}(\varepsilon)=&\{z\in\imath\mathbb{R}\colon\penalty 10000\ |\operatorname{Im}(z)|\geq\varepsilon\},\\[6.0pt] \Pi_{3}(\varepsilon)=&\{z\in\imath\mathbb{R}\colon\penalty 10000\ |\operatorname{Im}(z)|<\varepsilon\}.\end{split}$ (A.16)
@@
$\left|\frac{S_{2}(z)}{S_{2}(z+g\omega_{2})}\right|=\big|\big(2\sin(\pi z)\big)^{g}\big|(1+O(\omega_{2})),\penalty 10000\ \penalty 10000\ z\in\Pi_{1}(\varepsilon),\penalty 10000\ \penalty 10000\ 0\leq\omega_{2}\leq a(\varepsilon).$ (A.17)
@@
$\left|\frac{S_{2}(z)}{S_{2}(z+g\omega_{2})}\right|=\big|\big(2\sin(\pi z)\big)^{g}\big|(1+O(\omega_{2})),\penalty 10000\ \penalty 10000\ z\in\Pi_{2}(\varepsilon),\penalty 10000\ \penalty 10000\ 0\leq\omega_{2}\leq a(\varepsilon).$ (A.18)
@@
$\left|\frac{S_{2}(z)}{S_{2}(z+g\omega_{2})}\right|\leq C(\varepsilon),\penalty 10000\ \penalty 10000\ z\in\Pi_{3}(\varepsilon),\penalty 10000\ \penalty 10000\ 0\leq\omega_{2}\leq a(\varepsilon).$ (A.19)
@@
$\left|K^{R}(x|1,\omega_{2},g\omega_{2})\right|=\left|\frac{S_{2}(\imath x+\frac{1+\omega_{2}}{2}+\frac{g\omega_{2}}{2})}{S_{2}(\imath x+\frac{1+\omega_{2}}{2}-\frac{g\omega_{2}}{2})}\right|=\left|2\operatorname{ch}(\pi x)\right|^{-g}\left(1+O(\omega_{2})\right)\!.$ (A.20)
@@
$\begin{split}&\left|\mu^{R}(x|1,\omega_{2},g\omega_{2})\right|=\left|\frac{S_{2}(\imath x)}{S_{2}(\imath x+g\omega_{2})}\frac{S_{2}(-\imath x)}{S_{2}(-\imath x+g\omega_{2})}\right|=\left|4\operatorname{sh}^{2}(\pi x)\right|^{g}\left(1+O(\omega_{2})\right),\end{split}$ (A.21)
@@
$\log\frac{S_{2}(z+g\omega_{2})}{S_{2}(z)}=\frac{\pi\imath g}{2}\left({2z}+{(g-1)\omega_{2}}-1\right)+\!\!\int\limits_{\mathbb{R}+\imath 0}\!\!\frac{e^{zt}(e^{g\omega_{2}t}-1)}{(e^{t}-1)(e^{\omega_{2}t}-1)}\frac{dt}{t},$ (A.24)
@@
$\begin{split}&\log\frac{S_{2}(z)}{S_{2}(z+g\omega_{2})}-g\log\big(2\sin(\pi z)\big)=-\frac{\pi\imath\omega_{2}}{2}g(g-1)-\\[6.0pt] &\int_{\mathbb{R}+\imath 0}\frac{e^{zt}}{e^{t}-1}\left(\frac{e^{g\omega_{2}t}-1}{e^{\omega_{2}t}-1}-g\right)\frac{dt}{t}=-\,v.p.\int_{\mathbb{R}}\frac{e^{zt}}{e^{t}-1}\left(\frac{e^{g\omega_{2}t}-1}{e^{\omega_{2}t}-1}-g\right)\frac{dt}{t}\end{split}$ (A.25)
@@
$\log\big|2\sin(\pi z)\big|=\pi\operatorname{Im}z-\mathrm{Re}\,\int\limits_{\mathbb{R}+\imath 0}\frac{e^{zt}}{e^{t}-1}\frac{dt}{t},\,\quad 0\leq\mathrm{Re}\,(z)<1,\quad z\not=0,$ (A.26)
@@
$\hskip 113.81102pt0\leq\mathrm{Re}\,(z)<1+(1-g)\omega_{2},\quad z\not=0.$
@@
Proofs of Propositions A.1A.3. Using Lemmas A.1, A.2 and A.3 we obtain
$\begin{split}\left|\mathrm{Re}\,\left(\log\frac{S_{2}(z)}{S_{2}(z+g\omega_{2})}-g\log\big(2\sin(\pi z)\big)\right)\right|\leq\omega_{2}(C_{1}+C_{2}+C_{3}),\\ 0\leq\omega_{2}\leq a(\varepsilon),\penalty 10000\ \penalty 10000\ z\in\Pi_{1}(\varepsilon).\end{split}$
@@
$\begin{split}e^{-\omega_{2}(C_{1}+C_{2}+C_{3})}\leq\left|\frac{S_{2}(z)}{S_{2}(z+g\omega_{2})}\cdot\big(2\sin(\pi z)\big)^{-g}\right|\leq e^{\omega_{2}(C_{1}+C_{2}+C_{3})},\\ 0\leq\omega_{2}\leq a(\varepsilon),\penalty 10000\ \penalty 10000\ z\in\Pi_{1}(\varepsilon),\end{split}$
@@
To prove Proposition A.3 apply Lemmas A.1, A.2 and A.5 to (LABEL:closed)